A Complete Bibliography of Publications in ACM SIGSOFT Software Engineering Notes: 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 August 2018 Version 1.00

Title word cross-reference

1976 [Rin77]. 1979 [Bai79].

52779 [Sou78].

A. [DLP78]. abstract [Lin78b, Wür78]. Abstraction [LSAS77, SWL77, Nou78, Sof79d]. Abstracts [Sof76, Sof77a, Sof77b, Sof78a, Sof78b, Sof78c, Sof79a, Sof79b]. acceptance [Mar78]. ACM [Par77a]. actions [Lom77]. Ages [Dij78]. aid [Sta78]. algebraic [GHM77]. Alphard [SWL77]. alternative [Sof79c, YS78]. Analysis [Las78, WK77]. and/or [Sch79b]. annotated [Kle78]. Anthony [Rin77]. application [Sti78]. applications [SG78]. applicative [FW77]. Applied [Pfa78]. approach [BT79]. Approaches [Zav79, Bow78, Nou78]. April [Bai79]. areas [BT79]. Aspects [FW77]. Association [Par77a]. assurance [BB78, BS78, Fis78, Fuj78, Kac78, Pfa78, SG78, Tig78]. atomic [Lom77]. automatic [BFFG78].

Babb [Ano80]. based [Jos78]. be [Mey78]. before [Lam78]. better [Las78]. beyond [Mar78]. bibliography [Gil78b, Kle78]. BLOWHARD [Par77b]. book [Mar78]. Building [Par77b].

Carriage [Par77a]. charts [YS78]. checking [FL77]. CLU [LSAS77]. Code [Gla78b]. Comment [Gil78b, Gil79]. commentary [Van78]. communications [Goo77]. comparison [Fuj78, Gel79a, Par78a]. completeness [Lin78a]. components [Eme79]. Computer [TW78, Jos78, Mey78]. Computing [Gla78a]. conference [SH79, Bai79]. Configuration [BHS78]. Constructing [Goo77]. consultants [Sch79b]. Control [Gla78b, Jos78]. controversy [Par78b]. Correction [Ano80]. correctness [DW78, Mau78]. current [Bow78]. curriculum [Fai79a, Fai79b].

data [Buc77, Gan77, Lin78b, Nou78, Sti78]. Defining [SWL77, BT79]. definition [Gil79]. DeMillo [DLP78, Van78, Kil79]. describing [Lam78]. Design [Gel79a, Sta78, Gil78a, Kin77, Las78, Loc78, Par78a, Pet78, PHL^{+}77, RSSW77, Rob77, YS78]. designs [RW78, YW78]. detection [WK77]. determination [CC77]. development [BB78, DB78, Kac78, Rad78]. Dijkstra [Jef78, Kil79, Par78b, Van78]. direct [Gil78a]. discipline [Wir77]. dMLP [Par78b]. down [Was76]. DREAM [RSSW77]. dynamic [CC77].

Early [GMS77]. education [Rin77]. effect [Lov77]. effective [DR78]. effectiveness [BB78]. Efficient [FL77]. effort [BB78]. electronic [BFFG78]. enforcement [Buc77]. engineering [BT79, Fai79a, Fai79b, Kle78, Rin77, Was76, Zel78]. Error [SG78, Ott78, WK77]. establishment [YW78]. Euclid [PHL^{+}77]. evaluate [YW78]. evaluation [BB78, Gan77]. evolution [Gla78b]. example [Kac78]. Exception [Mac77, MSR77]. exchange [Ger78]. execution [BFFG78]. experience [Fis78, GMS77, Gla78a]. experienced [BB78]. experiences [Bai79]. experiment [BS78]. experimental [Gan77, Lov77]. extensions [GHM77]. failure [Gla78a]. Farshid [Lin78a]. features [AM77, BT77]. feedback [Sta78]. field [BT79]. file [FW77]. Florida [Ham79]. flowcharts [YS78]. formal [Lin78a]. FORTRAN [WK77]. framework [CM78]. Freeman [Rin77]. function [BFFG78]. Further [Ott78]. future [Fis78].

generalized [CC77]. generation [BFFG78]. generators [SWL77]. goals [Gil78a]. goes [Mar78]. graph [Sti78]. Gypsy [AGB^{+}77].

handling [Mac77, MSR77]. hardware [BT77]. High [BT77]. historical [Par77a]. hypothesis [Ott78].

IDBUG [DB78]. ideal [Her77]. ideas [Pri78]. IEEE [SH79]. implementation [AGB^{+}77, FL77]. Improvement [Woo78]. informal [Zav79]. interaction [AM77]. introduction [RSSW77]. invariant [Buc77]. investigation [Lov77]. item [Par77a]. iteration [SWL77].

J [DLP78].

Karl [Gil78b]. key [Las78]. Kleine [Gil78b].

Language [AM77, AGB^{+}77, BT77, DR78, Her77, Kin77]. Languages [RW78, AH77, Cob78]. large [BB78, Kac78]. large-scale [BB78]. learning [Gla78a]. Lebesgue [Sha78]. Letter [Gel79a, Kil79, Sha78, Van78, Mau78]. level [BT77]. library [Jos78]. limitations [Kin77]. Limits [SH79]. Lipton [DLP78, Kil79, Van78]. logic [Nou78, Sha78]. logic-oriented [Nou78].

maintainability [Gel79b, Gil79].

Management [BHS78, Rei76]. Managing [Rad78]. Manufacturers [Par77a]. Masters [Fai79a, Fai79b]. measurement [CM78, Cob78, Zel78]. measures [YW78]. measuring [Bow78]. mechanisms [LSAS77]. members [Sch79a]. Mesa [GMS77]. method [Mar79]. methodological [Har78]. methodologies [Gel79a, Par78a]. methodology [Pfa78, Pri78, Tig78, TW78]. methods [Fuj78]. metrics [Loc78]. Middle [Dij78]. MIL [Sou78]. MIL-S-52779 [Sou78]. mini [Jos78]. mini-computer [Jos78]. modelling [SG78]. models [DW78, Zav79]. Multidimensional [Gil78a].

Nassi [YS78]. needs [Rin77]. next [Ard79]. note [Har78, Hor79, Nou78, Sof79d]. Notes [PHL^{+}77]. notices [Par77a]. Nourani [Lin78a].

objectives [Rin77]. operating [Wür78]. optimization [FL77]. oriented [Nou78].

Pamphlet [Dij78]. Panel [Kin77, Zav79, Bai79, SH79]. paper [Ano80, Dij77, Jef78]. Parker [Gel79a]. Parnas [Kil79]. part [Sof77a, Sof77b, Sof78a, Sof78b, Sof78c, Sof79a, Sof79b]. PASCAL [FL77]. performance [Loc78]. Perlis [DLP78, Kil79, Van78]. Peter [Rin77]. PGN [Sof79d]. PL [Mac77]. PL/I [Mac77]. plans [BFFG78]. Political [Dij78]. position [Dij77, Jef78]. Preliminary [FW77]. probabilistic [DW78]. problem [Lam78, Sch79b]. Process [Lom77, AM77, Rob77]. processes [Gil78a]. processing [DR78, Goo77]. Productivity [Zel78]. Program [Woo78, DB78, DW78, Ger78, Hor79, Lov77, Sch79b, Sha78]. programmed [MSR77]. programming [AH77, Cob78, FW77, Gan77, Her77, Sof79c, TW78, Wir77]. programs [AGB^{+}77, Mar79, Mau78, Mey78]. progress [SH79]. projects [Zel78]. proofs [Ger78]. proper [Tig78]. properties [Buc77, CC77]. proposal [Ger78]. protection [AH77]. psychology [Sof79d]. publication [Ger78].

Quality [Woo78, BB78, BS78, Bow78, CM78, Fis78, Kac78, Pfa78, SG78, Tig78, YW78]. quantified [Gil78a]. quantitative [BB78].

R. [DLP78]. real [DR78, Mau78, Van78, Wir77]. real-time [DR78, Wir77]. recovery [Lom77]. references [Sch79b]. regarding [Van78]. Relating [Pet78]. reliability [Dij77, Gan77, Hor79, Jef78, MSR77]. Reliable [Bai79, Rob77, Ard79, Goo77, Kin77, Par77b, SH79]. remarks [Wür78]. reply [Lin78a, Van78]. Report [Ham79, Sch79a, Zav79]. representing [RW78]. request [Sch79b]. Requirements [DR78, Pie78, Fis78, Pet78]. Responding [Sou78]. Response [DLP78]. Restricted [Buc77]. restriction [Lin78b]. Review [Rin77, Har78, Pri78]. role [MSR77]. run [FL77]. run-time [FL77].

S [Sou78]. scale [BB78, Eme79, Kac78]. Secretary [Sch79a]. Secretary-Treasurer [Sch79a]. Selected [Kle78, Gil78b]. selection [Sti78]. service [Mil79]. session [Zav79]. Shneiderman [YS78]. Should [Mey78, Gil78a]. SIGSOFT [Sch79a]. SIGTRANS [Par77a]. Small [Eme79]. Small-scale [Eme79]. smart [Rei76]. Software [Bai79, BHS78, Fis78, KP76, Mar78, MSR77, Woo78, Ard79, BT79, BB78, BS78, Bow78, CM78, Dij77, Eme79, Fai79a, Fai79b, Fuj78, Gil78a, Jef78, Kac78, Kin77, Kle78, Las78, Mil79, Par77b, Pet78, Rad78, Rei76, RSSW77, RW78, Rob77, SH79, Sof76, Sof77a, Sof77b, Sof78a, Sof78b, Sof78c, Sof79a, Sof79b, Sta78, Sti78, SG78, Tig78, Van78, Was76, YW78, Zel78, Rin77]. solution [Lam78]. Some [GHM77, Mil79, Pri78, Wür78]. Source [Gla78b]. special [BT77]. specification [AGB^{+}77, Bai79, Buc77, Zav79]. specification-various [Zav79].

Specifications [Ard79, GHM77, Lin78a, RW78, SH79, Wür78, Bai79]. Specifying [Lin78b, Mar79, SWL77]. Springer [Rin77]. Springer-Verlag [Rin77]. State [Lam78]. Static [CC77, WK77]. statistics [Mil79]. structure [Cob78, Lov77]. structured [Mar79, Sof79c]. structuring [Lom77]. stub [Rei76]. study [AH77]. successful [Kac78]. sufficient [Bow78]. switching [BFFG78]. synchronization [Lom77]. syndicate [Sof79c]. System [Gla78b, Jos78, Loc78, RSSW77]. systems [BFFG78, DR78, FW77, Goo77, Sta78, Wür78].

techniques [Kac78]. test [BFFG78, Sti78]. Testing [Gel79b, DR78, Ham79, Mar78, Mil79, Sch79b, Sha78]. theory [Sti78]. time [DR78, FL77, Wir77]. Tool [Pie78, DB78, Rei76]. tools [Bai79, Fis78, KP76]. top [Was76]. top-down [Was76]. Tracing [Pie78]. Treasurer [Sch79a]. Tripp [Ano80]. type [CC77]. types [Buc77, Gan77, Lin78b].

understanding [Lov77]. unions [CC77]. unstructured [Cob78]. use [Sta78, Wür78, YW78]. used [Kac78]. User [Bai79]. Using [Loc78, Lom77]. utilization [BT77].

validation [Ott78]. value [Tig78]. variables [Buc77]. various [Zav79]. verifiable [AGB^{+}77]. verification [SWL77]. verified [Goo77, Mey78]. Verlag [Rin77]. Version [FW77]. view [Kil79, Par78b, Was76]. vs [Kil79].

Wasserman [Rin77]. within [BT79]. workshop [Ham79].

References

Ambler:1977:GLS

[AGB^{+}77] Allen L. Ambler, Donald I. Good, James C. Browne, Wilhelm F. Burger, Richard M. Cohen, Charles G. Hoch, and Robert E. Wells. Gypsy: a language for specification and implementation of verifiable programs. ACM SIGSOFT Software Engineering Notes, 2(2):1–10, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ambler:1977:SPP

[AH77] Allen L. Ambler and Charles G. Hoch. A study of protection in programming languages. ACM SIGSOFT Software Engineering Notes, 2(2):25–40, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Andrews:1977:LFP

[AM77] Gregory R. Andrews and James R. McGraw. Language features for process interaction. ACM SIGSOFT Software Engineering Notes, 2(2):114–127, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1980:CBT

[Ano80] Anonymous. Correction to Babb and Tripp paper. ACM SIGSOFT Software Engineering Notes, 5 (1):29, January 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [BT79].

Ardis:1979:SRS

[Ard79] Mark A. Ardis. Specifications of reliable software — what next? ACM SIGSOFT Software Engineering Notes, 4(3):19–21, July 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bail:1979:UES

[Bai79] William G. Bail. User experiences with specification tools: (panel from Specifications of Reliable Software Conference) (April 3–5, 1979). ACM SIGSOFT Software Engineering Notes, 4(3):7– 14, July 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Belford:1978:QEE

[BB78] Peter Chase Belford and Carlo Broglio. A quantitative evaluation of the effectiveness of quality assurance as experienced on a large-scale software development effort. ACM SIGSOFT Software Engineering Notes, 3(5):173– 180, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bauer:1978:AGE

[BFFG78] Jonathan Bauer, Susan Faasse, Alan Finger, and William Goodhue. The automatic generation and execution of function test plans for electronic switching systems. ACM SIGSOFT Software Engineering Notes, 3(5):92– 100, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bersoff:1978:SCM

[BHS78] Edward H. Bersoff, Vilas D. Henderson, and Stan G. Siegel. Software configuration management. ACM SIGSOFT Software Engineering Notes, 3(5):9–17, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowen:1978:CAS

[Bow78] John B. Bowen. Are current approaches sufficient for measuring software quality? ACM SIGSOFT Software Engineering Notes, 3(5): 148–155, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benson:1978:SQA

[BS78] J. P. Benson and S. H. Saib. A software quality assurance experiment. ACM SIGSOFT Software Engineering Notes, 3(5):87–91, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brown:1977:HLL

[BT77] R. Leonard Brown and Paul G. Tuttle. High level language utilization of special hardware features. ACM SIGSOFT Software Engineering Notes, 2(5):14–16, October 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Babb:1979:ADA

[BT79] Robert G. Babb and Leonard L. Tripp. An approach to defining areas within the field of software engineering. ACM SIGSOFT Software Engineering Notes, 4(4): 9–17, October 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See correction [Ano80].

Buckle:1977:RDT

[Buc77] Normand Buckle. Restricted data types, specification and enforcement of invariant properties of variables. ACM SIGSOFT Software Engineering Notes, 2(2): 68–76, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cousot:1977:SDD

[CC77] Patrick Cousot and Radhia Cousot. Static determination of dynamic properties of generalized type unions. ACM SIGSOFT Software Engineering Notes, 2(2): 77–94, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cavano:1978:FMS

[CM78] Joseph P. Cavano and James A.

McCall. A framework for the measurement of software quality. ACM SIGSOFT Software Engineering Notes, 3(5):133–139, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cobb:1978:MSU

[Cob78] Gary W. Cobb. A measurement of structure for unstructured programming languages. ACM SIGSOFT Software Engineering Notes, 3(5):140–147, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Drasch:1978:ITP

[DB78] Frederick J. Drasch and Richard A. Bowen. IDBUG: a tool for program development. ACM SIGSOFT Software Engineering Notes, 3(5): 106–110, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dijkstra:1977:PPSb

[Dij77] Edsger W. Dijkstra. A position paper on software reliability. ACM SIGSOFT Software Engineering Notes, 2(5):3–5, October 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dijkstra:1978:PPM

[Dij78] Edsger W. Dijkstra. On a political pamphlet from the Middle Ages. ACM SIGSOFT Software Engineering Notes, 3(2):14–16, April 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See responses [DLP78, Jef78, Van78, Mau78].

DeMillo:1978:RRD

[DLP78] Richard A. DeMillo, Richard J. Lipton, and Alan J. Perlis. Response from R. A. DeMillo, R. J. Lipton, A. J. Perlis. ACM SIGSOFT Software Engineering Notes, 3(2):16–17, April 1978. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78].

Davis:1978:RLP

[DR78] Alan M. Davis and Walter J.

Rataj. Requirements language processing for the effective testing of real-time systems. ACM SIGSOFT Software Engineering Notes, 3(5): 61–66, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Duran:1978:TMP

[DW78] Joe W. Duran and John J.

Wiorkowski. Toward models for probabilistic program correctness. ACM SIGSOFT Software Engineering Notes, 3(5):39–44, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Emery:1979:SSS

[Eme79] James E. Emery. Small-scale software components. ACM SIGSOFT Software Engineering Notes, 4(4): 18–21, October 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fairley:1979:MCSa

[Fai79a] Richard E. Fairley. Masters curriculum in software engineering. ACM SIGSOFT Software Engineering Notes, 4(1):12–17, January 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fairley:1979:MCSb

[Fai79b] Richard E. Fairley. Masters curriculum in software engineering. ACM SIGSOFT Software Engineering Notes, 4(2):13–16, April 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fischer:1978:SQA

[Fis78] Kurt F. Fischer. Software quality assurance tools: Recent experience and future requirements. ACM SIGSOFT Software Engineering Notes, 3(5):116–121, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fischer:1977:EIO

[FL77] Charles N. Fischer and Richard J. LeBlanc. Efficient implementation and optimization of runtime checking in PASCAL. ACM SIGSOFT Software Engineering Notes, 2(2):19–24, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fujii:1978:CSA

[Fuj78] Marilyn S. Fujii. A comparison of software assurance methods. ACM SIGSOFT Software Engineering Notes, 3(5):27–32, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Friedman:1977:AAP

[FW77] Daniel P. Friedman and David S. Wise. Aspects of applicative programming for file systems (preliminary version). ACM SIGSOFT Software Engineering Notes, 2(2): 41–55, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gannon:1977:EED

[Gan77] J. D. Gannon. An experimental evaluation of data types on programming reliability. ACM SIGSOFT Software Engineering Notes, 2(2):141, March 1977. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Geller:1979:LPD

[Gel79a] Dennis P. Geller. Letter on Parker’s “Design methodologies: a comparison”. ACM SIGSOFT Software Engineering Notes, 4 (1):18, January 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Par78a].

Gelperin:1979:TM

[Gel79b] David Gelperin. Testing maintainability. ACM SIGSOFT Software Engineering Notes, 4(2):7–12, April 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerhart:1978:PPE

[Ger78] Susan L. Gerhart. A proposal for publication and exchange of program proofs. ACM SIGSOFT Software Engineering Notes, 3(1):7–17, January 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guttag:1977:SEA

[GHM77] Jhon V. Guttag, Ellis Horowitz, and David R. Musser. Some extensions to algebraic specifications. ACM SIGSOFT Software Engineering Notes, 2(2):63–67, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilb:1978:MQG

[Gil78a] Tom Gilb. Multidimensional quantified goals should direct software design processes. ACM SIGSOFT Software Engineering Notes, 3 (3):26–28, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gilkey:1978:CKK

[Gil78b] Thomas Gilkey. Comment on Karl Kleine’s “selected bibliography”. ACM SIGSOFT Software Engineering Notes, 3(2):19, April 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Kle78].

Gilb:1979:CDM

[Gil79]  Tom Gilb.  A comment on

“The definition of maintainability”. ACM SIGSOFT Software Engineering Notes, 4(3):32–33, July 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:1978:CFL

[Gla78a] Robert L. Glass. Computing failure: a learning experience. ACM SIGSOFT Software Engineering Notes, 3(5):18–19, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glasser:1978:ESC

[Gla78b] Alan L. Glasser. The evolution of a source code control system. ACM SIGSOFT Software Engineering Notes, 3(5):122– 125, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Geschke:1977:EEM

[GMS77] Charles M. Geschke, James H.

Morris, Jr., and Edwin H. Satterthwaite. Early experience with Mesa. ACM SIGSOFT Software Engineering Notes, 2(2):138, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Good:1977:CVR

[Goo77] Donald I. Good. Constructing verified and reliable communications processing systems. ACM SIGSOFT Software Engineering Notes, 2(5):8–13, October 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hamlet:1979:RFT

[Ham79] Richard Hamlet. Report on Florida testing workshop. ACM SIGSOFT Software Engineering Notes, 4(2):17–18, April 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrell:1978:NMR

[Har78] Herbert Harrell. A note on methodological review. ACM SIGSOFT Software Engineering Notes, 3(3):5–7, July 1978. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Herriot:1977:TIP

[Her77] Robert G. Herriot. Towards the ideal programming language. ACM SIGSOFT Software Engineering Notes, 2(2):56–62, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horning:1979:NPR

[Hor79] Jim Horning. A note on program reliability. ACM SIGSOFT Software Engineering Notes, 4(4):6–8, October 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeffrey:1978:DPP

[Jef78] H. J. Jeffrey. On Dijkstra’s position paper on software reliability. ACM SIGSOFT Software Engineering Notes, 3(2):18, April 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78].

Josephs:1978:MCB

[Jos78] William H. Josephs. A minicomputer based library control system. ACM SIGSOFT Software Engineering Notes, 3(5):126– 132, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kacik:1978:ESQ

[Kac78] Paul J. Kacik. An example of software quality assurance techniques used in a successful large scale software development. ACM SIGSOFT Software Engineering Notes, 3(5):181–186, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kilov:1979:LPV

[Kil79] H. Kilov. Letter on Parnas’ view of Dijkstra vs. deMillo, Lipton and Perlis. ACM SIGSOFT Software Engineering Notes, 4(1):19, January 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

King:1977:PLL

[Kin77] James C. King. Panel: limitations of language design for reliable software. ACM SIGSOFT Software Engineering Notes, 2(3):11– 12, April 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kleine:1978:SAB

[Kle78] Karl Kleine. Selected annotated bibliography on software engineering. ACM SIGSOFT Software Engineering Notes, 3(1):18–25, January 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See comment [Gil78b].

Kernighan:1976:ST

[KP76] B. W. Kernighan and P. J.

Plauger. Software tools. ACM SIGSOFT Software Engineering Notes, 1(1):15–20, May 1976. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lamport:1978:SPB

[Lam78] Leslie Lamport. State the problem before describing the solution. ACM SIGSOFT Software Engineering Notes, 3(1):26, January 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lassagne:1978:AKB

[Las78] Theodore D. Lassagne. Analysis is the key to better software design. ACM SIGSOFT Software Engineering Notes, 3(3):24– 26, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Linden:1978:FSC

[Lin78a] Theodore A. Linden. On formal specifications and completeness: reply to Farshid Nourani. ACM SIGSOFT Software Engineering Notes, 3(3):15–17, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Nou78].

Linden:1978:SAD

[Lin78b] Theodore A. Linden. Specifying abstract data types by restriction. ACM SIGSOFT Software Engineering Notes, 3(2):7–13, April 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lockett:1978:UPM

[Loc78] Joann Lockett. Using performance metrics in system design. ACM SIGSOFT Software Engineering Notes, 3(5):156–159, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lomet:1977:PSS

[Lom77] D. B. Lomet. Process structuring, synchronization, and recovery using atomic actions. ACM SIGSOFT Software Engineering Notes, 2(2):128–137, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Love:1977:EIE

[Lov77] Tom Love. An experimental investigation of the effect of program structure on program understanding. ACM SIGSOFT Software Engineering Notes, 2(2):105–113, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liskov:1977:AMC

[LSAS77] Barbara Liskov, Alan Snyder, Russell Atkinson, and Craig Schaffert. Abstraction mechanisms in CLU. ACM SIGSOFT Software Engineering Notes, 2(2):140, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

MacLaren:1977:EHP

[Mac77] M. Donald MacLaren. Exception handling in PL/I. ACM SIGSOFT Software Engineering Notes, 2(2): 101–104, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Martin:1978:SAT

[Mar78] K. A. Martin. Software acceptance testing that goes beyond the book. ACM SIGSOFT Software Engineering Notes, 3(5):101– 105, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marca:1979:MSS

[Mar79] David Marca. A method for specifying structured programs. ACM SIGSOFT Software Engineering Notes, 4(3):22–31, July 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Maurer:1978:ALC

[Mau78] W. D. Maurer. Another letter on the correctness of real programs. ACM SIGSOFT Software Engineering Notes, 3(3):22–24, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78, DLP78, Jef78, Van78].

Meyer:1978:SCP

[Mey78] Steve Meyer. Should computer programs be verified? ACM SIGSOFT Software Engineering Notes, 3(3):18–19, July 1978. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miller:1979:SSS

[Mil79] Edward F. Miller. Some statistics from the software testing service. ACM SIGSOFT Software Engineering Notes, 4(1):8–11, January 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Melliar-Smith:1977:SRR

[MSR77] P. M. Melliar-Smith and B. Randell. Software reliability: The role of programmed exception handling. ACM SIGSOFT Software Engineering Notes, 2(2):95–100, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nourani:1978:NLO

[Nou78] Farshid Nourani. A note on logicoriented approaches to data abstraction. ACM SIGSOFT Software Engineering Notes, 3(3):13– 15, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See reply [Lin78a].

Ottenstein:1978:FVE

[Ott78] Linda M. Ottenstein. Further validation of an error hypothesis. ACM SIGSOFT Software Engineering Notes, 3(1):27–28, January 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1977:AAC

[Par77a] Dave Parnas. ACM (association of carriage manufacturers): SIG-TRANS notices — another historical item. ACM SIGSOFT Software Engineering Notes, 2(5):6–7, October 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:1977:BRS

[Par77b] David L. Parnas. Building reliable software in BLOWHARD. ACM SIGSOFT Software Engineering Notes, 2(3):5–6, April 1977. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parker:1978:CDM

[Par78a] John Parker. A comparison of design methodologies. ACM SIGSOFT Software Engineering Notes, 3(4):12–19, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See letter [Gel79a].

Parnas:1978:AVD

[Par78b] David L. Parnas. Another view of the Dijkstra–dMLP controversy. ACM SIGSOFT Software Engineering Notes, 3(4):20–21, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Peters:1978:RSR

[Pet78] Lawrence Peters. Relating software requirements and design. ACM SIGSOFT Software Engineering Notes, 3(5):67–71, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfau:1978:AQA

[Pfa78] Pamela R. Pfau. Applied quality assurance methodology. ACM SIGSOFT Software Engineering Notes, 3(5):1–8, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Popek:1977:NDE

[PHL^{+}77] G. J. Popek, J. J. Horning, B. W. Lampson, J. G. Mitchell, and R. L.

London. Notes on the design of Euclid. ACM SIGSOFT Software Engineering Notes, 2(2):11–18, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pierce:1978:RTT

[Pie78] Robert A. Pierce. A requirements tracing tool. ACM SIGSOFT Software Engineering Notes, 3(5): 53–60, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Priebe:1978:SIR

[Pri78]  Kathie Priebe.  Some ideas

on review methodology. ACM SIGSOFT Software Engineering Notes, 3(4):27–29, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raduchel:1978:MSD

[Rad78] William J. Raduchel. Managing software development. ACM SIGSOFT Software Engineering Notes, 3(4):22–26, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reifer:1976:SSS

[Rei76] Donald J. Reifer. The smart stub as a software management tool. ACM SIGSOFT Software Engineering Notes, 1(2):4–8, October 1976. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rine:1977:RSE

[Rin77] David Rine. Review of Software engineering education: needs and objectives by Anthony I. Wasserman and Peter Freeman. Springer-Verlag 1976. ACM SIGSOFT Software Engineering Notes, 2(3):32, April 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robinson:1977:RSD

[Rob77] L. Robinson. Reliable software and the design process. ACM SIGSOFT Software Engineering Notes, 2(3):7–10, April 1977. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riddle:1977:IDS

[RSSW77] William E. Riddle, John H. Sayler, Alan R. Segal, and Jack C. Wileden. An introduction to the DREAM software design system. ACM SIGSOFT Software Engineering Notes, 2(4):11–24, July 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Riddle:1978:LRS

[RW78] William E. Riddle and Jack C. Wileden. Languages for representing software specifications and designs. ACM SIGSOFT Software Engineering Notes, 3(4):7–11, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schneidewind:1979:STR

[Sch79a] Norm Schneidewind. Secretary-Treasurer’s report to SIGSOFT members. ACM SIGSOFT Software Engineering Notes, 4(3):3– 5, July 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schorer:1979:PTP

[Sch79b] Pete Schorer. A program testing problem: request for references and/or consultants. ACM SIGSOFT Software Engineering Notes, 4(4):22, October 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sukert:1978:EMA

[SG78] Alan N. Sukert and Amrit L.

Goel. Error modelling applications in software quality assurance. ACM SIGSOFT Software Engineering Notes, 3(5):33–38, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shore:1979:LSW

[SH79] John E. Shore and Kathryn L.

Heninger. Limits to specifications: why not more progress? Panel at IEEE conference on specifications of reliable software. ACM SIGSOFT Software Engineering Notes, 4(3):15–16, July 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaw:1978:LLL

[Sha78] Mary Shaw. Letter on Lebesgue, logic, and program testing. ACM SIGSOFT Software Engineering Notes, 3(4):21, October 1978. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1976:ASE

[Sof76] Software Engineering Notes Staff. Abstracts in software engineering. ACM SIGSOFT Software Engineering Notes, 1(2):9–32, October 1976. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1977:ASEa

[Sof77a] Software Engineering Notes Staff. Abstracts in software engineering: part 2. ACM SIGSOFT Software Engineering Notes, 2(1):6–14, January 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1977:ASEb

[Sof77b] Software Engineering Notes Staff. Abstracts in software engineering: part 3. ACM SIGSOFT Software Engineering Notes, 2(4):24– 28, July 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1978:ASEa

[Sof78a] Software Engineering Notes Staff. Abstracts in software engineering: part 5. ACM SIGSOFT Software Engineering Notes, 3(1): 28–32, January 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1978:ASEb

[Sof78b] Software Engineering Notes Staff. Abstracts in software engineering: part 6. ACM SIGSOFT Software Engineering Notes, 3(3):31– 33, July 1978. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1978:ASEc

[Sof78c] Software Engineering Notes Staff.

Abstracts in software engineering: part 6. ACM SIGSOFT Software Engineering Notes, 3(4):31–33, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1979:ASEa

[Sof79a] Software Engineering Notes Staff.

Abstracts in software engineering: part 8. ACM SIGSOFT Software Engineering Notes, 4(1): 22–24, January 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1979:ASEb

[Sof79b] Software Engineering Notes Staff.

Abstracts in software engineering: part 9. ACM SIGSOFT Software Engineering Notes, 4(2):19– 20, April 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1979:ASP

[Sof79c] Software Engineering Notes Staff.

An alternative to structured programming: syndicate programming. ACM SIGSOFT Software Engineering Notes, 4(1):19– 20, January 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staff:1979:NPA

[Sof79d] Software Engineering Notes Staff.

A note on the psychology of abstraction (PGN). ACM SIGSOFT Software Engineering Notes, 4 (1):21, January 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Southworth:1978:RM

[Sou78] Richard N. Southworth. Responding to MIL-S-52779. ACM SIGSOFT Software Engineering Notes, 3(5):160–164, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stavely:1978:DFU

[Sta78] Allan M. Stavely. Design feedback and its use in software design aid systems. ACM SIGSOFT Software Engineering Notes, 3(5): 72–78, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stickney:1978:AGT

[Sti78] M. E. Stickney. An application of graph theory to software test data selection. ACM SIGSOFT Software Engineering Notes, 3(5): 111–115, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaw:1977:AVA

[SWL77] Mary Shaw, Wm A. Wulf, and Ralph L. London. Abstraction and verification in Alphard: Defining and specifying iteration and generators. ACM SIGSOFT Software Engineering Notes, 2(2):139, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tighe:1978:VPS

[Tig78] Michael F. Tighe. The value of a proper software quality assurance methodology. ACM SIGSOFT Software Engineering Notes, 3(5): 165–172, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Turski:1978:CPM

[TW78] W. M. Turski and Anthony I.

Wasserman. Computer programming methodology. ACM SIGSOFT Software Engineering Notes, 3(2):20–21, April 1978. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

VanGhent:1978:LRS

[Van78] Roger Van Ghent. Letter on real software, regarding the commentary by Dijkstra and the reply by DeMillo, Lipton and Perlis. ACM SIGSOFT Software Engineering Notes, 3(3):20–21, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78, DLP78, Jef78].

Wasserman:1976:TVS

[Was76] Anthony Ira Wasserman. A topdown view of software engineering. ACM SIGSOFT Software Engineering Notes, 1(1):8–14, May 1976. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wirth:1977:TDR

[Wir77] N. Wirth. Towards a discipline of real-time programming. ACM SIGSOFT Software Engineering Notes, 2(2):142, March 1977. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wendel:1977:FED

[WK77] Irv. K. Wendel and Richard L. Kleir. FORTRAN error detection through static analysis. ACM SIGSOFT Software Engineering Notes, 2(3):22–28, April 1977. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Woodmancy:1978:SQI

[Woo78] Donald A. Woodmancy. A software quality improvement program. ACM SIGSOFT Software Engineering Notes, 3(5):20– 26, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wurges:1978:SRU

[Wür78] Harald Würges. Some remarks on the use of abstract specifications for operating systems. ACM SIGSOFT Software Engineering Notes, 3(3):8–12, July 1978. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yoder:1978:NSC

[YS78] Cornelia M. Yoder and Marilyn L. Schrag. Nassi–Shneiderman charts an alternative to flowcharts for design. ACM SIGSOFT Software Engineering Notes, 3(5):79– 86, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yin:1978:EUM

[YW78] B. H. Yin and J. W. Winchester.

The establishment and use of measures to evaluate the quality of software designs. ACM SIGSOFT Software Engineering Notes, 3(5): 45–52, November 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zave:1979:PSA

[Zav79] Pamela Zave. Panel session: Approaches to specification-various models: an informal report. ACM SIGSOFT Software Engineering Notes, 4(3):17–18, July 1979. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zelkowitz:1978:PMS

[Zel78] Marvin V. Zelkowitz. Productivity measurement on software engineering projects. ACM SIGSOFT Software Engineering Notes, 3(4): 30–31, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).